# StructLM: Towards Building Generalist Models for Structured Knowledge Grounding

### v2

@@ 5 Ablation Studies
Figure 4: Effect of different pretraining curricula on SKG finetuning performance in relevant task groupings.
@@
Figure 5: Effect of general instruction-following data on held-out SKG dataset performance. Performance is measured as the average over evaluation metrics across all tasks within held-in or held-out groups. Note that the held-in performance experiences a milder dip compared to the held-out performance gains.
Effect of base model pretraining data. We ablate our choice of base model, CodeLlama-7b-Instruct, by finetuning the unspecialized Llama2-7b base model and Llemma, which is further pretrained on mathematical texts Azerbayev et al. (2023b). Intuitively, one might guess that coding ability has the most transferability to performance on the types of SKG tasks we are studying due to the symbolic nature of programming languages and code-writing scenarios. However, it is possible that other types of pretraining to boost reasoning ability, such as math, have even greater transferability.
@@
Effect of general instruction data in mixture We ablate our inclusion of approximately $500$k general instruction-following examples into the training data mixture to observe the importance of including this data. As we see in Figure 5, the held-in performance is relatively unaffected by the added general examples, but held-out performance improves significantly with more general data. Empirically, we also observe that when training a large volume of task-specific input and output formats, the model becomes less capable of following instructions on new tasks in a zero-shot setting. We hypothesize that training on this general mixture helps zero-shot performance because it can reduce overfitting to the task formats in the training set.
Cross-task and cross-format transferability We ablate the transferability of performance between input structure knowledge types and between output task types. To test this, we train a number of tasks together and compare them to their single-task finetuned models. Our results are indicative that there is cross-task transferability of performance occurring. In schema and knowledge triples, training on different tasks seems to weakly improve the performance of the evaluation task compared to finetuning on only that evaluation task. On tables, we see that this effect is much greater. One explanation for this is that there are a wide variety of table tasks included in the training mix, and the volume of auxiliary data is larger than the other data types.
On the other hand, we see that finetuning on different datatypes with the same task (i.e. summarization) also yields benefits to performance. On the summarization and question-answering (QA) experiments, we train on both tabular and knowledge graph data. We evaluate summarization with Dart and QA with WikiSQL. We see that even though we have only added one extra dataset in each case, we still see an improved result compared to the single-task model even though these datasets are have very different knowledge types. Thus, overall evidence supports that the performance of tasks within our training set are benefiting from each other, even across data types.
@@ 6 Discussion
We argue that SKG is an important capability for future language models. We have seen through our experiments on ChatGPT and the Llama2 family that there is significant room for improvement.
We found that we could produce a strong model by focused instruction-tuning on SKG tasks, however, we also observe that the performance difference between 7B to 34B StructLM models was not dramatic. This raises a concern about the state of SKG data: could we be approaching a performance ceiling? Combined with the fact that we were able to outperform UL2-20b, a much larger model, with our 7B model on 3 tasks, it seems that LLMs at various scales are struggling with SKG capabilities.
@@ 7 Conclusion
In this paper, we explore the current capabilities of open language models on structured knowledge grounding tasks. We show that LLMs are currently weak at SKG tasks currently. To address this gap, we construct an instruction-tuning dataset mixture of 1.1M examples and release models that outperform USKG on a majority ($14$ of $18$) tasks and achieve SOTA on $7$ of them. We also study the effects of various factors that influence the performance of our model on these task types. We hope that our work provides an updated understanding of what is achievable in the SKG domain, and can serve as a strong baseline for future improvements.
Limitations
Additionally, while we have tried to evaluate our models to the best of our ability, many of the tasks of our held-out datasets measure accuracy through a heuristic matching step of a model’s output. In zero or few-shot settings, it is quite challenging to exactly control the generations of an autoregressive transformer to be adherent to a certain rule or grammar, and this has been a subject of study in other works (i.e. PICARD Scholak et al. (2021)). We note that because of this reality, poor results in zero or few-shot context may betray the existence of useful representations that the model has already learned. Without further prompting or finetuning efforts, it may be difficult to bring these capabilities to light. As such, another opportunity for improvement upon our methods may involve more flexible constrained methods of language model evaluation.
Ethics Statement
This paper does not raise any ethical concerns. The data and additional resources employed in this study are open-source and widely utilized in existing works.
Appendix A Appendix
@@
GrailQA EM val 77.1 TIARA + GAIN (T5-3B) (Shu and Yu, 2023)
@@
Infotabs Acc dev 75.6 Infotabs paper(Gupta et al., 2020a)
@@
Table 4: Performance comparison of various models on different datasets, highlighting the SOTA scores and their sources. When selecting SoTA scores, we exclude methods that use large closed-source models such as GPT4 with agent-based methods. We also specify the data split that this score was achieved on.
@@
Table 5: Ablation results for the mixtures of general data in the training set. In total, we train 5 models, where the percentage represents the percent of the training data that is general. In the held out data, we see noticeable gains in generalization performance for FinQA and InfoTabs datasets. Notably, FinQA requires the generation of a python-executable math expression and InfoTabs requires an exact match to 3 previously unseen (boolean) options. WikiTableText performance seems to suffer, but is evaluated based the BLEU score with only one target sentence. As a result, we place more emphasis on the model’s 0-shot adaptation ability to new output specifications unseen in the training data.
@@
Table 6: The evaluation results comparing finetuning done on different base models. Code refers to CodeLlama-Instruct-7B. Math refers to Llemma-7b. LLaMA refers to Llama2-7b.

### v3

@@ 5 Ablation Studies
Figure 4: Effect of different pretraining curricula on SKG finetuning performance in relevant task groupings. We can observe the advantages of CodeLlma over the others.
@@
Figure 5: Effect of general instruction-following data on averaged held-out SKG dataset performance. Performance is measured as the average over evaluation metrics across all tasks within held-in or held-out groups.
Effect of base model pretraining data. We ablate our choice of base model, CodeLlama-7b-Instruct, by finetuning the unspecialized Llama2-7b base model and Llemma, which is further pretrained on mathematical texts (Azerbayev et al., 2023b). Intuitively, one might guess that coding ability has the most transferability to performance on the types of SKG tasks we are studying due to the symbolic nature of programming languages and code-writing scenarios. However, it is possible that other types of pretraining to boost reasoning ability, such as math, have even greater transferability.
@@
Effect of general instruction data in mixture As we see in Figure 5, the held-in performance is relatively unaffected by the added general examples, but held-out performance improves significantly with more general data. Empirically, we also observe that when training a large volume of task-specific input and output formats, the model becomes less capable of following instructions on new tasks in a zero-shot setting. We hypothesize that training on this general mixture helps zero-shot performance because it can reduce overfitting to the task formats in the training set.
Cross-task and cross-format transferability We ablate the transferability of performance between input structure knowledge types and between output task types. To test this, we train a number of tasks together and compare them to their single-task finetuned models. Our results are indicative that there is cross-task transferability of performance occurring. On tables, we see this effect clearly. This may be explained by the volume and variety of table tasks included in the training mix. In schema and knowledge triples, the performance improvement is not pronounced, perhaps due to the limited size of those datasets. Nevertheless, evidence supports that diversifying tasks on the same structured knowledge type benefits performance.
On the other hand, we see that finetuning on different datatypes with the same task (i.e. summarization) also yields benefits to performance. On the summarization and question-answering (QA) experiments, we train on both tabular and knowledge graph data. We evaluate summarization with DART and QA with WikiSQL. We see that in both cases, the extra dataset yielded about 1% improvement. Considering that the added datasets in each case organize information in a completely different way, this result suggests diversifying data types for similar tasks do indeed benefit each other as well.
@@ 6 Discussion
We argue that SKG is an important capability for future language models. We have seen through our experiments on ChatGPT and the Llama2 family that there is significant room for improvement. We found that we could produce a strong model by focused instruction-tuning on SKG tasks, however, we also observe that the performance difference between 7B to 34B StructLM models was not dramatic. This raises a concern about the state of SKG data: could we be approaching a performance ceiling? Combined with the fact that we were able to outperform UL2-20b, a much larger model, with our 7B model on 3 tasks, it seems that LLMs at various scales are struggling with SKG capabilities.
@@ 7 Conclusion
In this paper, we explore the current capabilities of open language models on structured knowledge grounding tasks. We show that LLMs are currently weak at SKG tasks. To address this gap, we construct an instruction-tuning dataset mixture of 1.1M examples and release models that and achieve SOTA on $7$ of 18 held-in tasks, and that outperform strong existing models such as TableLlama and Flan UL2 on held-out tasks. We also study the effects of various factors that influence the performance of our model on these task types. We hope that our work provides an updated understanding of what is achievable in the SKG domain, and can serve as a strong baseline for future improvements.
Appendix A SoTA Results
@@
GrailQA EM val 77.1 TIARA + GAIN (T5-3B) (Shu & Yu, 2023)
@@
Infotabs Acc dev 75.6 Infotabs paper(Gupta et al., 2020b)
@@
Table 4: Specification of SoTA scores and their sources on the most prevalent metrics used for assessment.
Across the datasets that we evaluate, the SoTA scores are chosen based on methods that do not use agent based methods on models as large as GPT-4 for a fairer comparison. Across these datasets, we can see that many of the best performing methods are purpose-built for the type of data structure used in the task. For example, RESDSQL focuses exclusively on controlled SQL generation, DATER uses SQL-based reasoning for tabular tasks, and S3HQA focuses on table and text multi-hop QA. Compared to these methods, the performance of StructLM can be seen a strong baseline to determine if these domain-specific designs yield real benefits.
Appendix B SlimOrca Dataset Mixture Details
@@
Table 5: Ablation results for the mixtures of general data in the training set.
In total, we train 5 models, where the percentage represents the percent of the training data that is general. In the held out data, we see noticeable gains in generalization performance for FinQA and InfoTabs datasets. Notably, FinQA requires the generation of a python-executable math expression and InfoTabs requires an exact match to 3 previously unseen (boolean) options. WikiTableText performance seems to suffer, but is evaluated based on the BLEU score with only one target sentence. As a result, we place more emphasis on the model’s 0-shot adaptation ability to new output specifications unseen in the training data.
Appendix C Per-dataset Pretraining Data Comparison
@@
Table 6: Fine-grained evaluation results comparing finetuning done on different base models. Code refers to CodeLlama-Instruct-7B. Math refers to Llemma-7b. LLaMA refers to Llama2-7b.
In these fine-grained results we can compare the performance of Llemma, Llama, and CodeLlama in more detail. Llemma does seem to hold an advantage on tasks that involve math (TabMWP, FinQA). Math pretraining does not seem to benefit tabular tasks overall (WikiSQL, WikiTQ, HybridQA, MMQA, etc.), but does show advantages on SQL coding tasks (Spider, SparC). CodeLlama, however, seems to show a performance improvement over the base Llama model on not just coding tasks, but also math and tabular tasks as well. These findings may underscore the need to understand what constitutes ”reasoning” in a language model.
Appendix D Prompt Format
@@
Appendix E Held-Out Generation Examples
For illustration purposes, we provide examples of successful and unsuccessful responses of StructLM-13B on the FinQA held-out dataset.
E.1 Successful StructLM-13B Response on FinQA
Use the financial data in the following passage and table to calculate the answer the question. table: performance graph the following graph compares the yearly change in the cumulative total stockholder return for our last five full fiscal years , based upon the market price of our common stock , with the cumulative total return on a nasdaq composite index ( u.s . companies ) and a peer group , the nasdaq medical equipment-sic code 3840-3849 index , which is comprised of medical equipment companies , for that period . the performance graph assumes the investment of \$ 100 on march 31 , 2007 in our common stock , the nasdaq composite index ( u.s . companies ) and the peer group index , and the reinvestment of any and all dividends. . col : |3/31/2007 |3/31/2008 |3/31/2009 |3/31/2010 |3/31/2011 |3/31/2012 row 1 : abiomed inc |100 |96.19 |35.87 |75.55 |106.37 |162.45 row 2 : nasdaq composite index |100 |94.11 |63.12 |99.02 |114.84 |127.66 row 3 : nasdaq medical equipment sic code 3840- |100 |82.91 |41.56 |77.93 |94.54 |74.40 this graph is not 201csoliciting material 201d under regulation 14a or 14c of the rules promulgated under the securities exchange act of 1934 , is not deemed filed with the securities and exchange commission and is not to be incorporated by reference in any of our filings under the securities act of 1933 , as amended , or the exchange act whether made before or after the date hereof and irrespective of any general incorporation language in any such filing . transfer agent american stock transfer & trust company , 59 maiden lane , new york , ny 10038 , is our stock transfer agent. . question: what is the roi of an investment in nasdaq composite index from march 2007 to march 2010? write down a short python expression.   (99.02 - 100) / 100
In the above example, StructLM-13B identified the relevant quantity within the given linearized table, then determined how to calculate the required ROI quantity in a python-executable expression.
E.2 Unsuccessful StructLM-13B Response on FinQA
Use the financial data in the following passage and table to calculate the answer the question. table: performance graph the following graph and table compares the cumulative five-year total return provided to shareholders on our class a common stock relative to the cumulative total returns of the s&p 500 index and our customized peer group . the peer group includes cboe holdings , inc. , intercontinentalexchange group , inc . and nasdaq , inc . an investment of \$ 100 ( with reinvestment of all dividends ) is assumed to have been made in our class a common stock , in the peer group and the s&p 500 index on december 31 , 2012 , and its relative performance is tracked through december 31 , 2017 . comparison of 5 year cumulative total return* among cme group inc. , the s&p 500 index , and a peer group 12/12 12/13 12/14 12/15 12/16 cme group inc . s&p 500 peer group * \$ 100 invested on 12/31/12 in stock or index , including reinvestment of dividends . fiscal year ending december 31 . copyright a9 2018 standard & poor 2019s , a division of s&p global . all rights reserved . the stock price performance included in this graph is not necessarily indicative of future stock price performance. . col : |2013 |2014 |2015 |2016 |2017 row 1 : cme group inc . |\$ 164.01 |\$ 194.06 |\$ 208.95 |\$ 279.85 |\$ 370.32 row 2 : s&p 500 |132.39 |150.51 |152.59 |170.84 |208.14 row 3 : peer group |176.61 |187.48 |219.99 |249.31 |323.23 unregistered sales of equity securities during the past three years there have not been any unregistered sales by the company of equity securities. . question: in 2017 what was the ratio of the the cme group inc . stock perfomamce to the s&p write down a short python expression.   (cme_group_inc_stock_performance / s&p_500_stock_performance)
In the above example, StructLM-13B understood how to calculate the required quantity, but could not select the relevant numbers from the given table.
Limitations
Additionally, while we have tried to evaluate our models to the best of our ability, many of the tasks of our held-out datasets measure accuracy through a heuristic matching step of a model’s output. In zero or few-shot settings, it is quite challenging to exactly control the generations of an autoregressive transformer to be adherent to a certain rule or grammar, and this has been a subject of study in may of the other works cited in Table 4. We note that because of this reality, poor results in zero or few-shot context may betray the existence of useful representations that the model has already learned. Without further prompting or finetuning efforts, it may be difficult to bring these capabilities to light. As such, another opportunity for improvement upon our methods may involve more flexible constrained methods of language model evaluation.